OpenStudy (anonymous):

x=(15/y)+2 Please isolate y

OpenStudy (anonymous):

\[x=\frac{ 15 }{ y }+2\]

OpenStudy (tkhunny):

Have you considered subtracting 2?

OpenStudy (anonymous):

Yeah. I went through how I thought and ended up with 15 over x over 2

OpenStudy (tkhunny):

Hmmm... \(x = \dfrac{15}{y} + 2\) Subtract 2 \(x - 2 = \dfrac{15}{y} \) Reciprocals \(\dfrac{1}{x - 2} = \dfrac{y}{15} \) Multiply by 15 \(\dfrac{15}{x - 2} = y \) Not quite what you said, is it?

OpenStudy (anonymous):

No but I looked back in my notes and figured it out